CKET TU CLASS QUESTIO ORDER: 30 MG/M2 HEIGHT: 92 CM WEIGHT: 17 KG AVAILABLE: 5MG/ML HOW MANY ML WILL YOU GIVE?
Answer #1

the order is 30mg /M2

height is 92 cm that is 0.92 m

so the person should have the m2 value will be 0.922 = 0.8464

so inorder to give 30mg / m2 the person should recieve 30× 0.8464 = 23.36064 mg

available dose is 5mg per ml

so 1mg = 1/5 ml = 0.2ml

so 23.6064 mg = 0.2× 23.36064 ml = 4.672128ml